OpenStudy (idealist10):

Find a particular solution of y"-y'+y=e^(x)(2+x)sinx.

OpenStudy (idealist10):

@freckles

OpenStudy (freckles):

look at the factors on the right hand side and look at the table I gave you

OpenStudy (freckles):

\[y''-y'+y=e^{x} \text{ you might guess } y=Ae^{x} \\ y''-y'+y =2+x \text{ you might guess } y=Ax+B \\ y''-y'+y=\sin(x) \text{ you might guess } y=A \sin(x)+B \cos(x)\]

OpenStudy (idealist10):

@zepdrix

OpenStudy (idealist10):

@pooja195 @Luigi0210 @Kainui

OpenStudy (daniel.ohearn1):

Ay, solve for y" then substitute after solving for y prime, y prime in terms of y should should allow you to isolate y in terms of x. From there it should be more apparent.

OpenStudy (freckles):

what is the question here?